OpenStudy (maheshmeghwal9):

Solve: (8-a)(8-b)(8-c)(8-d)(8-e)=245.Then the value of {a+b+c+d+e} will be equal to? a,b,c,d,e are distinct positive integers.

OpenStudy (blockcolder):

Are a, b, c, d, e positive integers?

OpenStudy (maheshmeghwal9):

yes!

OpenStudy (anonymous):

Must be, otherwise there will be more than two answers

OpenStudy (blockcolder):

245=5*7^2. WLOG, a=3, b=c=1, d=e=7. a+b+c+d+e=3+2+14=19

OpenStudy (anonymous):

Factorize 245, you'll get 5.7.7. Assign first three (8-x)'s to these and the rest to 1. So 8-5,8-7,8-7,8-1,8-1 are your numbers.

OpenStudy (maheshmeghwal9):

ya

OpenStudy (maheshmeghwal9):

nooo nooooo

OpenStudy (maheshmeghwal9):

8-3,8-7,8-7,8-1,8-1

OpenStudy (ajprincess):

It is an interesting question.

OpenStudy (maheshmeghwal9):

ya

OpenStudy (maheshmeghwal9):

but there are 4 options are given for this question 5,18,24,35.But issacdean gave answer 19

OpenStudy (maheshmeghwal9):

pleas you solve.

OpenStudy (blockcolder):

It's supposed to be 19.

OpenStudy (maheshmeghwal9):

no.

OpenStudy (maheshmeghwal9):

but thanx for your co-operation. Now I will solve it

OpenStudy (ajprincess):

Bt I too get the same answr I mean 19.

OpenStudy (maheshmeghwal9):

ok np.:( .I Think answer given is wrong.

OpenStudy (ajprincess):

I tried several times bt getting the same answer :(.

OpenStudy (maheshmeghwal9):

ok let it go.Thanx!

OpenStudy (ajprincess):

yw.

OpenStudy (maheshmeghwal9):

bye